OpenStudy (anonymous):

prove cos x + sin x tan x= 1/cosx

OpenStudy (lgbasallote):

take the left hand side... \[\cos x + \sin x \tan x\] change tanx to sinx/cosx \[\implies \cos x + \sin x ( \frac{\sin x}{\cos x})\] simplify \[\implies \cos x + \frac{\sin^2 x}{\cos x}\] change to common denominators \[\implies \frac{\cos^2 x}{\cos x} + \frac{\sin^2 x}{\cos x}\] add \[\implies \frac{\cos^2 x + \sin^2 x}{\cos x}\] now.. you know that cos^2 x + sin^2 x = 1 because of pythagorean identity \[\implies \frac 1{\cos x} = RHS\] does that help?